OpenStudy (anonymous):

HELP ME! A child’s wading pool has a diameter of 5 feet and a height of 1 foot. How much water would it take to fill the pool? Round your answer to the nearest gallon. (Hint: 1 cubic foot of water is approximately 7.5 gallons.) 79 gallons 147 gallons 59 gallons 589 gallons

OpenStudy (anonymous):

well I am going to assume this child's wading pool is a cylinder because that is all that comes to mind at the moment. If this is true, the formula is V=pi(r)^2h pi as in 3.14...... the diameter is 5 feet so to find the radius divide it by 2 therefore r=2.5 V=pi(2.5)^2(1) V=19.63 (about) cubic feet then multiple this by 7.5 for gallons